X Company plans to prepare annual financial statements. After the company's first year of operation, the following financial statement information is available:
Balance Sheet
Total assets $15,561
Total liabilities 6,065
Total Paid-In Capital 4,165
Income Statement
Total expenses $17,603
What was total revenue for the year?
A: $14,319 B: $16,754 C: $19,602 D: $22,934 E: $26,833 F: $31,394 G: $36,731 H: $42,976
Answer #1
Total assets = Total liabilities + Total shareholders equity
Total assets = Total liabilities + (Total paid in capital + Retained earnings)
Retained earnings = Net Income = (Total Revenue - Total Expenses)
15561 = 6065+(4165+(Total Revenue - 17603))
(Total Revenue - 17603) = (15561-6065-4165)
(Total Revenue - 17603) = 5331
Total Revenue = 17603+5331
Total Revenue = $ 22,934/.
Option D) is the answer